OpenStudy (anonymous):

________ dividing without using long division: 6 ) 4598 how do you do it without using division or multiplication

OpenStudy (asnaseer):

well, one way is to keep subtracting 6 from the number and count how many times you did that

OpenStudy (anonymous):

what do you mean by that.. can you like show it to me.. once please thanks

OpenStudy (asnaseer):

lets say you wanted to divide 18 by 6, then the steps would be as follows: 1. 18 - 6 = 12 2. 12 - 6 = 6 3. 6 - 6 = 0 (finished) so answer is 3 because we subtracted 6 a total of 3 times

OpenStudy (anonymous):

Thankyou!! :)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

wait.. one more thing.. so the answer is 766 for this question.. do i have to like subtract it 766.. wouldn't it take like forever lol

OpenStudy (asnaseer):

it would - which is why we usually use long division or a calculator. :)

OpenStudy (asnaseer):

there is also a remainder of 2 in this problem

OpenStudy (anonymous):

yeah cuz im doing my summer assignment and it strictly says not to use calculator, long division or multiplication.. soo im soo confused.. yeah i know that

OpenStudy (asnaseer):

you could also use grouping

OpenStudy (anonymous):

wait. how doyou do that? lol.. i knw how but hwo to do it in this one..

OpenStudy (asnaseer):

so, for example, add 6 to itself 10 times and you will get 60 then see how many times you can subtract 60 from 4598

OpenStudy (anonymous):

76 r 2

OpenStudy (asnaseer):

or even this: 1. add 6 to itself 10 times to get 60 2. add 60 to itself 10 times to get 600 3. add 600 to itself 7 times to get 4200

OpenStudy (asnaseer):

then use these groups to subtract bigger chucks out of your number until you can no longer subtract anything else

OpenStudy (asnaseer):

BTW: 766 r2

OpenStudy (asnaseer):

do you see how creating a larger group will reduce the number of operatons you need to do?

OpenStudy (anonymous):

yes... thankyou soo much ...welll i have like alot of more questions..

OpenStudy (asnaseer):

yw. <--- and please post each question one at a time in the list to your left

OpenStudy (anonymous):

oh okay.. thanks

OpenStudy (asnaseer):

np :)

OpenStudy (asnaseer):

you need to click on the blue Close button before you can ask another question. :)

OpenStudy (asnaseer):

look at the bottom right hand side of your question area if you cannot find it.